### v1

# Quench dynamics via recursion method and dynamical quantum phase transitions

Abstract
The recursion method provides a powerful framework for studying quantum many-body dynamics in the Lanczos basis, which is recursively constructed within the Krylov space of operators. Recently, it has been demonstrated that the recursion method, when supplemented by the universal operator growth hypothesis, can effectively compute autocorrelation functions and transport coefficients at infinite temperature. In this work, we extend the scope of the recursion method to far-from-equilibrium quench dynamics. We apply the method to spin systems in one, two, and three spatial dimensions. In one dimension we find a clear evidence that the convergence of the method is limited in time by the dynamical quantum phase transition.
Introduction.  Addressing quantum dynamics by means of the recursion method has a long history Mori (1965); Dupuis (1967); Haydock (1980); Mattis (1981); Joslin and Gray (1986); Roldan et al. (1986); Viswanath and Müller (2008); Lindner and Auerbach (2010); Nandy et al. (2024). In simple words, the (Heisenberg version of the) method amounts to constructing an orthonormal Lanczos basis within the Krylov space of an evolving quantum observable and harnessing the simplifications this basis brings about. A new impact has been given to the method by the universal operator growth hypothesis Parker et al. (2019), which has informed about the generic pattern of the evolution operator in the Lanczos basis.
Up to now the practical application of the recursion method to quantum many-body systems has been limited almost exclusively by computing infinite-temperature autocorrelation functions and transport coefficients Parker et al. (2019); Khait et al. (2016); de Souza et al. (2020); Yates et al. (2020a, b); Yuan et al. (2021); Wang et al. (2024); Bartsch et al. (2024); Uskov and Lychkovskiy (2024); Teretenkov et al. (2024); Bhattacharyya et al. (2024). There the method has demonstrated high efficiency. This efficiency is particularly notable in application to two-dimensional systems Uskov and Lychkovskiy (2024); Teretenkov et al. (2024); Bhattacharyya et al. (2024), where traditional numerical methods often struggle. Another notable advantage of the recursion method is that it allows a straightforward symbolic implementation, when a single computation covers all parameters of the Hamiltonian Uskov and Lychkovskiy (2024); Teretenkov et al. (2024).
Formally, the output of the recursion method is the Heisenberg operator of an observable. This suggests that the scope of the method can be much broader than infinite-temperature correlation functions and should cover, in particular, the out-of-equilibrium dynamics after quantum quenches. Addressing the later in systems of dimension higher than one is known to be highly challenging, with limited body of results obtained to date Schmitt and Heyl (2020); Dziarmaga (2022); Bernier and Agarwal (2025); Begušić and Chan (2024); Loizeau et al. (2025a). Given the success of the recursion method in computing two-dimensional autocorrelation function, it seems tempting to try to apply it to quench dynamics.
Here we undertake exactly this attempt. The paper is organized as follows. We start from briefly introducing the recursion method and specifying it for addressing quench dynamics. Then we apply it to a paradigmatic one-dimensional mixed-field Ising model. The results are benchmarked by essentially exact numerical simulations. We demonstrate that the recursion method has a convergence domain limited in time by the dynamical quantum phase transition. Then some results for the transverse-field Ising model on the square and cubic lattices are presented. We conclude by discussing implications of our results and open questions posed by them.
Recursion method for quench dynamics.  Consider an observable given by a self-adjoint Shrödinger operator $A$. The same observable in the Heisenberg representation reads $A_{t}=e^{itH}A\,e^{-itH}$, where $H$ is the Hamiltonian of the system. The Heisenberg operator $A_{t}$ evolves within the Krylov space spanned by operators $A,{\mathcal{L}}A,{\mathcal{L}}^{2}A,\dots$, where ${\mathcal{L}}\equiv[H,\bullet]$ is the superoperator referred to as Liouvillian. $A_{t}$ satisfies the Heisenberg equation of motion $\partial_{t}A_{t}=i\,{\mathcal{L}}A_{t}$. Within the recursion method Viswanath and Müller (2008), this equation is addressed in the orthonormal Lanczos basis $\{O^{0},O^{1},O^{2},\dots\}$ constructed as follows: $O^{0}=A/\|A\|$, $A^{1}={\mathcal{L}}O^{0}$, $b_{1}=\|A^{1}\|$, $O^{1}=b_{1}^{-1}\,A^{1}$ and $A^{n}={\cal L}\,O^{n-1}-\,b_{n-1}O^{n-2}$, $b_{n}=\|A^{n}\|$, $O^{n}=b_{n}^{-1}A^{n}$ for $n\geq 2$. Here a scalar product in the space of operators is introduced according to $\big{(}A|B\big{)}\equiv\operatorname{\mbox{tr}}\big{(}A^{\dagger}B\big{)}/d$, where $d$ is the Hilbert space dimension (which is assumed to be finite). The scalar product entails the norm $\|A\|=\sqrt{(A|A)}$. The coefficients $b_{n}$ are referred to as Lanczos coefficients. In this basis the Liouvillian has the tridiagonal form,
${\mathcal{L}}=\begin{pmatrix}0&b_{1}&0&0&\dots\\[3.00003pt] b_{1}&0&b_{2}&0&\dots\\[3.00003pt] 0&b_{2}&0&b_{3}&\dots\\[3.00003pt] 0&0&b_{3}&0&\dots\\[3.00003pt] \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}.$ (1)
The solution of the Heisenberg equation for the observable $A_{t}$ can be expressed as
$A_{t}=\sum_{n=0}^{\infty}\phi^{n}_{t}\,O^{n},$ (2)
where $\phi^{n}_{t}\equiv(O^{n}|A_{t})$ are the solution of the system of coupled equations, which can be written in the matrix form as
$\partial_{t}\phi_{t}=$ $i\,{\mathcal{L}}\phi_{t}.$ (3)
Here $\phi_{t}=\|\phi^{n}_{t}\|_{n=0,1,2,\dots}$ is a column vector and ${\mathcal{L}}$ is the matrix (1).
In practice, one can explicitly compute only a finite number $n_{\rm max}$ of Lanczos coefficients. The rest of the coefficients should be extrapolated. The universal operator growth hypothesis Parker et al. (2019) (see also precursor work Liu and Müller (1990); Florencio et al. (1992); Zobov and Lundin (2006); Elsayed et al. (2014); Bouch (2015)) states that in a generic quantum many-body system Lanczos coefficients $b_{n}$ scale asymptotically linearly with $n$ at large $n$ (with a logarithmic correction for one-dimensional systems). This asymptotic behavior has been confirmed in various many-body models Cao (2021); Noh (2021); Heveling et al. (2022); Uskov and Lychkovskiy (2024); De et al. (2024). As a rule, the leading asymptotics can be easily extracted from the first dozen of Lanczos coefficients. However, extrapolating Lanczos coefficients by a linear function is a too rough approximation: the qualitative behavior of the solutions of the Heisenberg equation strongly depends on the subleading asymptotic terms Viswanath et al. (1994); Parker et al. (2019); Yates et al. (2020a, b); Dymarsky and Smolkin (2021); Bhattacharjee et al. (2022); Avdoshkin et al. (2024); Camargo et al. (2023); Uskov and Lychkovskiy (2024); Dodelson (2025). Guided by the results of ref. Gam , we employ the following extrapolation formula:
$b_{n}$ $\simeq\frac{\alpha\,n}{\log n}+\gamma+\frac{(-1)^{n}\alpha}{\log n+c}$ ${\rm for}~{}~{}1{\rm D},$ (4)
$b_{n}$ $\simeq\alpha\,n+\gamma+\frac{(-1)^{n}\alpha}{\log n+c}$ ${\rm for}~{}~{}2,3{\rm D}.$ (5)
The last alternating term is required whenever $\varphi^{0}_{t}$ relaxes to a nonzero value at $t\rightarrow\infty$ Gam , which is generally the case. Parameters $\alpha,\,\gamma$ and $c$ are found by fitting $n_{\rm max}$ explicitly computed Lanczos coefficients.
The knowledge of the Heisenberg operator $A_{t}$ of the observable allows one to address the quench dynamics, i.e. compute the time-dependent value $\langle A_{t}\rangle\equiv\operatorname{\mbox{tr}}\rho_{0}A_{t}$ of the observable that evolves after the system is initialized in an out-of-equilibrium state $\rho_{0}$. Within our approach, we compute $\langle A_{t}\rangle$ according to
$\langle A_{t}\rangle\simeq\sum_{n=0}^{n_{\rm max}}\phi^{n}_{t}\,\langle O^{n}\rangle,$ (6)
where $\langle O^{n}\rangle\equiv\operatorname{\mbox{tr}}\rho_{0}O^{n}$. We compute $O^{n}$ and $b_{n}$ for $n$ up to $n_{\rm max}$ symbolically directly in the thermodynamic limit, as described in Uskov and Lychkovskiy (2024). $\langle O^{n}\rangle$ is also computed symbolically. As for functions $\phi^{n}_{t}$, they are obtained by numerically solving coupled linear differential equations (3) with the tridiagonal matrix ${\mathcal{L}}$ truncated at some large $k$ (we use $k=1000$ in the examples below). The entries of this matrix for $n>n_{\rm max}$ are obtained from extrapolation of the $n_{\rm max}$ known Lanczos coefficients according to extrapolation formulae (4), (5).
Note that $\|A\|\,\phi^{0}_{t}=d^{-1}\,\operatorname{\mbox{tr}}A\,A_{t}$ is the infinite-temperature autocorrelation function of the observable $A$. Recently the recursion method supplemented by the UOGH has proven effective in computing such autocorrelation functions Parker et al. (2019); Wang et al. (2024); Uskov and Lychkovskiy (2024); Teretenkov et al. (2024). The out-of-equilibrium post-quench expectation value $\langle A_{t}\rangle$ is, however, a more complex object. Below we explore it in specific models.
Figure 1: Time evolution of the (a) Loschmidt rate and (b)–(d) total polarization along $z$ direction in the 1D Ising model (7) with $h_{x}=h_{z}=1$. The initial state is a product translation-invariant state (9) with the polarization $(p_{x},p_{y},p_{z})$ indicated on each plot. Orange line – essentially exact results of the numerical time evolution (see text for details). Blue line - results of the recursion method. The Loschmidt rate features multiple cusp singularities manifesting dynamical quantum phase transitions. The results of the recursion method depart from the exact numerical results at time $t_{*}$ close to the time $t_{c}$ of the second cusp (red dashed vertical line).
Figure 2: Scaling of the estimated breakdown time $t_{*}$ with respect to the number $n_{\rm max}$ of explicitly know Lanczos coefficients. Circles, squares and diamonds correspond to pure initial states polarized in $x$, $y$ and $z$ direction, respectively. The dashed horizontal line marks the time $t_{c}$ of the dynamical quantum phase transition. Clearly, $t_{*}$ exhibits a tendency to saturate around $t_{c}$.
Quench in one dimension.  We focus on a paradigmatic nonintegrable one-dimensional spin-$1/2$ chain – the mixed-field Ising model. Its Hamiltonian reads
$H=\sum\nolimits_{j}\sigma^{x}_{j}\sigma^{x}_{j+1}+h_{z}\sum\nolimits_{j}\sigma^{z}_{j}+h_{x}\sum\nolimits_{j}\sigma^{x}_{j},$ (7)
where $\sigma^{x,y,z}_{j}$ are Pauli matrices at the $j$’th site and $h_{x}$, $h_{z}$ are two parameters of the Hamiltonian. This model is integrable when $h_{x}=0$ or $h_{z}=0$, and nonintegrable otherwise.
The observable we consider here and throughout the paper is the magnetization in the $z$-direction,
$A=\sum\nolimits_{j}\sigma^{z}_{j}.$ (8)
Throughout the paper we consider a family of product translation-invariant initial states with polarization ${\bf p}$,
$\rho_{0}=\bigotimes_{j}\frac{1}{2}(\mathbb{1}+{\bf p}\bm{\sigma}_{j}),$ (9)
where ${\bf p}\bm{\sigma}_{j}=p_{x}\sigma^{x}_{j}+p_{y}\sigma^{y}_{j}+p_{z}\sigma^{z}_{j}$. In general, $|{\bf p}|\leq 1$; in particular, $|{\bf p}|=1$ for pure states with $\rho_{0}=|\Psi_{0}\rangle\langle\Psi_{0}|$.
We are able to compute $n_{\rm max}=45$ nested commutators and Lanczos coefficients. The time evolution of $\langle A_{t}\rangle$ computed by means of the recursion method is presented in Fig. 1. It is benchmarked by the results of exact diagonalization (performed with QuSpin Weinberg and Bukov (2017)) and time-dependent variational principle empowered by matrix product states (performed with ITensor Fishman et al. (2022a, b)). In one dimension the latter methods are essentially exact (and, of course, produce the same result) up to large times, therefore we regard their outcome as actual $\langle A_{t}\rangle$.
One can see from Fig. 1 that the recursion method excellently reproduces the actual $\langle A_{t}\rangle$ at initial time, but then abruptly breaks down. We observe that this breakdown is associated with a dynamical quantum phase transition (DQPT) Pollmann et al. (2010); Heyl et al. (2013); Karrasch and Schuricht (2013); Heyl (2018) – a phenomenon by which the Loschmidt rate function $\lambda(t)=N^{-1}\,\log\big{|}\langle\Psi_{0}|e^{-i\,t\,H}\Psi_{0}\rangle\big{|}^{2}$ features singularities in the thermodynamic limit. Namely, the breakdown time $t_{*}$ is very close tho the time $t_{c}$ of one of the singularities (specifically, the second singularity), see Fig. 1.
The breakdown time $t_{*}$ grows with the number $n_{\rm max}$ of available explicit Lanczos coefficient. However, this growth is saturated at $t_{c}$, as shown in Fig. 2. This strengthens our belief that the observed association between the breakdown of the recursion method and the DQPT is not accidental.
Remarkably, while the recursion method breaks down at a finite time $t_{*}$, at large times it seems to typically converge to the actual equilibrium value of the observable, as illustrated in Fig. 3 (a). Note that the one-dimensional Ising model is known for its atypical thermalization properties Bañuls et al. (2011); Lin and Motrunich (2017); Peng and Cui (2022); Pathak (2024), including long-lived Lin and Motrunich (2017) or even eternal Bañuls et al. (2011) oscillations of polarization around its time-averaged value seen in Fig. 3 (a).
Figure 3: Quench dynamics on (a) one-dimensional nearest-neighbor, (b) square, and (c) cubic lattices. Shown is the expectation value of the total magnetization in the $z$-direction. The initial state is the pure product state (9) polarized along the same direction. Blue solid line is the result of the recursion method. The chosen time span is sufficient for the latter to equilibrate. (a) One-dimensional Ising model (7) with $h_{x}=h_{z}=1$ (the same as in Fig. 1 but for a larger time span). Orange is the exact numerical result. The horizontal dashdotted line is the numerically time-averaged post-quench value of the observable. (b) and (c): The transverse-filed Ising model (10) with $h_{z}=1$ on the square and cubic lattices, respectively. Points are the results of the sparse Pauli dynamics method Begušić and Chan (2024).
Quench in higher dimensions.  In two and three dimensions we consider the transverse-field Ising model on the square and cubic lattices, respectively. The Hamiltonian of the transverse-field Ising model can be written as
$H=\sum_{\langle{\mathbf{i}}{\mathbf{j}}\rangle}\sigma^{x}_{{\mathbf{i}}}\sigma^{x}_{{\mathbf{j}}}+h_{z}\sum_{{\mathbf{j}}}\sigma^{z}_{{\mathbf{j}}},$ (10)
where ${\mathbf{i}}$ and ${\mathbf{j}}$ enumerate lattice sites and the first sum runs over pairs of neighbouring sites.
We compute $22$ and $12$ nested commutators, respectively. Some results are presented in Fig. 3 (b),(c). They are compared to the outcomes of the sparse Pauli dynamics method Begušić and Chan (2024); Begušić .
One can see that the agreement is good initially but abruptly worsens at some threshold time. We are not able to decisively resolve the origin of this discrepancy at the moment. Of course, the DQPT is again the main suspect. DQPT is known to occur in the 2D Ising model, although with a softer logarithmic singularity Heyl (2018). However, we are not able to compute the critical time for this DQPT for two- and three-dimensional models.
Conclusions.  In this work, we have explored the applicability of the recursion method to computing the time evolution of observables relaxing from out-of-equilibrium initial states. We have applied the method to quantum spin systems in one, two, and three dimensions. In one dimension, benchmarking is straightforward due to the availability of essentially exact numerical simulations. By comparing the recursion method to these simulations, we observed that it yields accurate results up to a certain threshold time, beyond which it abruptly breaks down. Notably, we found that this breakdown time closely coincides with the critical time of a dynamical quantum phase transition (DQPT), suggesting that DQPT fundamentally limits the applicability of the recursion method, at least in one dimension.
The underlying mechanism linking DQPT to the breakdown of the recursion method remains entirely unknown and represents a pressing open question. The potential subtlety of this mechanism is highlighted by our observation that the breakdown time coincides not with the first, but with the second (in time) dynamical critical point, implying that not every DQPT necessarily disrupts the recursion method.
In contrast to the one-dimensional case, benchmark data in two and three dimensions is scarce. We have compared our results with those obtained from the sparse Pauli dynamics method Begušić and Chan (2024); Begušić and found that the two approaches diverge abruptly after a certain threshold time, similar to what is observed in one dimension. However, due to the lack of available data on DQPTs in the higher-dimensional models we considered, we cannot unambiguously attribute this divergence to DQPTs. Determining whether DQPTs fundamentally constrain the recursion method’s convergence in higher dimensions remains an open question.
Finally, in one dimension, we have observed that at long times, the recursion method’s outcome coincides with the time-averaged value of the observable, which, in thermalizing systems, is equal to the equilibrium value. If this observation proves to be correct and generic, it would imply an unusual scenario: the recursion method remains valid at both short and long times but fails at intermediate times. Investigating this phenomenon further presents another promising direction for future research.
Note added.  While this manuscript has been in preparation, a closely related study has appeared Loizeau et al. (2025b). The latter paper addresses quench dynamics in one-dimensional spin systems by means of the recursion method. It does not report neither the finite convergence domain of the method nor an association with the DQPT.
Acknowledgements.
Acknowledgments.  We sincerely thank Alexey Rubtsov for pointing to the dynamical phase transition as a potential cause of the recursion method breakdown. We are grateful to Tomislav Begušić and Garnet Kin-Lic Chan for discussing their paper Begušić and Chan (2024), and particularly for providing unpublished data Begušić on the 2D Ising model. Numerous discussions with Alexander Teretenkov and Nikolay Il’in are acknowledged. This work was supported by the Russian Science Foundation under grant № 24-22-00331, https://rscf.ru/en/project/24-22-00331/

### v2

@@
 Abstract
-The recursion method provides a powerful framework for studying quantum many-body dynamics in the Lanczos basis, which is recursively constructed within the Krylov space of operators.
+The recursion method provides a powerful framework for studying quantum many-body dynamics in the Lanczos basis recursively constructed within the Krylov space of operators.
 Recently, it has been demonstrated that the recursion method, when supplemented by the universal operator growth hypothesis, can effectively compute autocorrelation functions and transport coefficients at infinite temperature.
-In this work, we extend the scope of the recursion method to far-from-equilibrium quench dynamics.
+We extend the scope of the recursion method to far-from-equilibrium quench dynamics.
 We apply the method to spin systems in one, two, and three spatial dimensions.
-In one dimension we find a clear evidence that the convergence of the method is limited in time by the dynamical quantum phase transition.
+In one dimension we find an evidence that the convergence of the method is limited in time by the dynamical quantum phase transition.
 Introduction.
@@
 (2024).
-In simple words, the (Heisenberg version of the) method amounts to constructing an orthonormal Lanczos basis within the Krylov space of an evolving quantum observable and harnessing the simplifications this basis brings about.
+The (Heisenberg version of the) method essentially amounts to constructing an orthonormal Lanczos basis within the Krylov space of an evolving quantum observable and harnessing the simplifications this basis brings about.
 A new impact has been given to the method by the universal operator growth hypothesis Parker et al.
-(2019), which has informed about the generic pattern of the evolution operator in the Lanczos basis.
-Up to now the practical application of the recursion method to quantum many-body systems has been limited almost exclusively by computing infinite-temperature autocorrelation functions and transport coefficients Parker et al.
+(2019) – a conjecture about a generic form of the evolution generator in the Lanczos basis.
+Up to now the practical application of the recursion method to quantum many-body systems has been limited almost exclusively to computing infinite-temperature autocorrelation functions and transport coefficients Parker et al.
 (2019); Khait et al.
@@
 (2024); Bhattacharyya et al.
+(2024); Füllgraf et al.
 (2025).
@@
 This suggests that the scope of the method can be much broader than infinite-temperature correlation functions and should cover, in particular, the out-of-equilibrium dynamics after quantum quenches.
-Addressing the later in systems of dimension higher than one is known to be highly challenging, with limited body of results obtained to date Schmitt and Heyl (2020); Dziarmaga (2022); Bernier and Agarwal (2025); Begušić and Chan (2024); Loizeau et al.
-(2025a).
-Given the success of the recursion method in computing two-dimensional autocorrelation function, it seems tempting to try to apply it to quench dynamics.
+Addressing the later in systems of dimension higher than one is known to be highly challenging, with limited body of results obtained to date Schmitt and Heyl (2020); Li et al.
+(2020); Dziarmaga (2022); Bernier and Agarwal (2025); Begušić and Chan (2025); Loizeau et al.
+(2025a, b).
+Given the success of the recursion method in computing two-dimensional autocorrelation function, it seems tempting to apply it to quench dynamics.
 Here we undertake exactly this attempt.
@@
 We demonstrate that the recursion method has a convergence domain limited in time by the dynamical quantum phase transition.
-Then some results for the transverse-field Ising model on the square and cubic lattices are presented.
-We conclude by discussing implications of our results and open questions posed by them.
-Recursion method for quench dynamics.
-Consider an observable given by a self-adjoint Shrödinger operator $A$.
-The same observable in the Heisenberg representation reads $A_{t}=e^{itH}A\,e^{-itH}$, where $H$ is the Hamiltonian of the system.
-The Heisenberg operator $A_{t}$ evolves within the Krylov space spanned by operators $A,{\mathcal{L}}A,{\mathcal{L}}^{2}A,\dots$, where ${\mathcal{L}}\equiv[H,\bullet]$ is the superoperator referred to as Liouvillian.
-$A_{t}$ satisfies the Heisenberg equation of motion $\partial_{t}A_{t}=i\,{\mathcal{L}}A_{t}$.
-Within the recursion method Viswanath and Müller (2008), this equation is addressed in the orthonormal Lanczos basis $\{O^{0},O^{1},O^{2},\dots\}$ constructed as follows: $O^{0}=A/\|A\|$, $A^{1}={\mathcal{L}}O^{0}$, $b_{1}=\|A^{1}\|$, $O^{1}=b_{1}^{-1}\,A^{1}$ and $A^{n}={\cal L}\,O^{n-1}-\,b_{n-1}O^{n-2}$, $b_{n}=\|A^{n}\|$, $O^{n}=b_{n}^{-1}A^{n}$ for $n\geq 2$.
-Here a scalar product in the space of operators is introduced according to $\big{(}A|B\big{)}\equiv\operatorname{\mbox{tr}}\big{(}A^{\dagger}B\big{)}/d$, where $d$ is the Hilbert space dimension (which is assumed to be finite).
-The scalar product entails the norm $\|A\|=\sqrt{(A|A)}$.
-The coefficients $b_{n}$ are referred to as Lanczos coefficients.
-In this basis the Liouvillian has the tridiagonal form,
+Then results for the transverse-field Ising model on the square and cubic lattices are presented.
+We conclude by discussing implications of our results and open questions raised.
+[…]
 ${\mathcal{L}}=\begin{pmatrix}0&b_{1}&0&0&\dots\\[3.00003pt] b_{1}&0&b_{2}&0&\dots\\[3.00003pt] 0&b_{2}&0&b_{3}&\dots\\[3.00003pt] 0&0&b_{3}&0&\dots\\[3.00003pt] \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}.$ (1)
 The solution of the Heisenberg equation for the observable $A_{t}$ can be expressed as
-$A_{t}=\sum_{n=0}^{\infty}\phi^{n}_{t}\,O^{n},$ (2)
-where $\phi^{n}_{t}\equiv(O^{n}|A_{t})$ are the solution of the system of coupled equations, which can be written in the matrix form as
+$A_{t}=\|A\|\,\sum_{n=0}^{\infty}\phi^{n}_{t}\,O^{n},$ (2)
+where $\phi^{n}_{t}\equiv(O^{n}|A_{t})/\|A\|$ are the solution of the system of coupled equations, which can be written in the matrix form as
 $\partial_{t}\phi_{t}=$ $i\,{\mathcal{L}}\phi_{t}.$ (3)
@@
 (2022); Uskov and Lychkovskiy (2024); De et al.
-(2024).
+(2024); Ermakov (2025).
 As a rule, the leading asymptotics can be easily extracted from the first dozen of Lanczos coefficients.
@@
 (2024); Camargo et al.
-(2023); Uskov and Lychkovskiy (2024); Dodelson (2025).
+(2023); Uskov and Lychkovskiy (2024); Dodelson (2025); Gamayun et al.
+(2025).
 Guided by the results of ref.
-Gam , we employ the following extrapolation formula:
+Gamayun et al.
+(2025), we employ the following extrapolation formula:
 $b_{n}$ $\simeq\frac{\alpha\,n}{\log n}+\gamma+\frac{(-1)^{n}\alpha}{\log n+c}$ ${\rm for}~{}~{}1{\rm D},$ (4)
 $b_{n}$ $\simeq\alpha\,n+\gamma+\frac{(-1)^{n}\alpha}{\log n+c}$ ${\rm for}~{}~{}2,3{\rm D}.$ (5)
-The last alternating term is required whenever $\varphi^{0}_{t}$ relaxes to a nonzero value at $t\rightarrow\infty$ Gam , which is generally the case.
+The last alternating term is required whenever $\varphi^{0}_{t}$ relaxes to a nonzero value at $t\rightarrow\infty$ Gamayun et al.
+(2025), which is generally the case.
 Parameters $\alpha,\,\gamma$ and $c$ are found by fitting $n_{\rm max}$ explicitly computed Lanczos coefficients.
@@
 The entries of this matrix for $n>n_{\rm max}$ are obtained from extrapolation of the $n_{\rm max}$ known Lanczos coefficients according to extrapolation formulae (4), (5).
-Note that $\|A\|\,\phi^{0}_{t}=d^{-1}\,\operatorname{\mbox{tr}}A\,A_{t}$ is the infinite-temperature autocorrelation function of the observable $A$.
+It should be emphasized that the sequence $\langle O^{n}\rangle$ generally does not have any apparent structure and is not expected to have any particular asymptotic behavior, in contrast to the sequence of Lanczos coefficients $b_{n}$.
+We further discuss this fact and illustrate it by examples in the Supplement sup .
+Note that $\phi^{0}_{t}=({\cal N}/\|A\|)\,\operatorname{\mbox{tr}}A\,A_{t}$ is the infinite-temperature autocorrelation function of the observable $A$, up to a constant factor.
 Recently the recursion method supplemented by the UOGH has proven effective in computing such autocorrelation functions Parker et al.
@@
 The results of the recursion method depart from the exact numerical results at time $t_{*}$ close to the time $t_{c}$ of the second cusp (red dashed vertical line).
-Figure 2: Scaling of the estimated breakdown time $t_{*}$ with respect to the number $n_{\rm max}$ of explicitly know Lanczos coefficients.
-Circles, squares and diamonds correspond to pure initial states polarized in $x$, $y$ and $z$ direction, respectively.
-The dashed horizontal line marks the time $t_{c}$ of the dynamical quantum phase transition.
+Figure 2: Scaling of the estimated breakdown time $t_{*}$ with the number $n_{\rm max}$ of explicitly know Lanczos coefficients for various initial polarizations.
+$t_{*}$ is defined as the time when the recursion method result departs from the essentially exact numerical result by 0.005.
+The dashed horizontal line marks the time $t_{c}$ of the dynamical quantum phase transition, see Fig.
+1 (a).
 Clearly, $t_{*}$ exhibits a tendency to saturate around $t_{c}$.
@@
 In general, $|{\bf p}|\leq 1$; in particular, $|{\bf p}|=1$ for pure states with $\rho_{0}=|\Psi_{0}\rangle\langle\Psi_{0}|$.
-We are able to compute $n_{\rm max}=45$ nested commutators and Lanczos coefficients.
-The time evolution of $\langle A_{t}\rangle$ computed by means of the recursion method is presented in Fig.
+We are able to compute $n_{\rm max}=48$ nested commutators and Lanczos coefficients for the Hamiltonian (7) and the observable (8).
+The time evolution of $\langle A_{t}\rangle$ computed by means of the recursion method, i.e.
+from eq.
+(6), is presented in Fig.
 1.
-It is benchmarked by the results of exact diagonalization (performed with QuSpin Weinberg and Bukov (2017)) and time-dependent variational principle empowered by matrix product states (performed with ITensor Fishman et al.
-(2022a, b)).
-In one dimension the latter methods are essentially exact (and, of course, produce the same result) up to large times, therefore we regard their outcome as actual $\langle A_{t}\rangle$.
-One can see from Fig.
-1 that the recursion method excellently reproduces the actual $\langle A_{t}\rangle$ at initial time, but then abruptly breaks down.
-We observe that this breakdown is associated with a dynamical quantum phase transition (DQPT) Pollmann et al.
-(2010); Heyl et al.
-(2013); Karrasch and Schuricht (2013); Heyl (2018) – a phenomenon by which the Loschmidt rate function $\lambda(t)=N^{-1}\,\log\big{|}\langle\Psi_{0}|e^{-i\,t\,H}\Psi_{0}\rangle\big{|}^{2}$ features singularities in the thermodynamic limit.
-Namely, the breakdown time $t_{*}$ is very close tho the time $t_{c}$ of one of the singularities (specifically, the second singularity), see Fig.
-1.
-The breakdown time $t_{*}$ grows with the number $n_{\rm max}$ of available explicit Lanczos coefficient.
-However, this growth is saturated at $t_{c}$, as shown in Fig.
-2.
-This strengthens our belief that the observed association between the breakdown of the recursion method and the DQPT is not accidental.
-Remarkably, while the recursion method breaks down at a finite time $t_{*}$, at large times it seems to typically converge to the actual equilibrium value of the observable, as illustrated in Fig.
-3 (a).
-Note that the one-dimensional Ising model is known for its atypical thermalization properties Bañuls et al.
-(2011); Lin and Motrunich (2017); Peng and Cui (2022); Pathak (2024), including long-lived Lin and Motrunich (2017) or even eternal Bañuls et al.
-(2011) oscillations of polarization around its time-averaged value seen in Fig.
-3 (a).
-Figure 3: Quench dynamics on (a) one-dimensional nearest-neighbor, (b) square, and (c) cubic lattices.
+Values of $b_{n}$, as well as individual ingredients entering eq.
+(6), i.e.
+$\phi^{n}_{t}$ and $\langle O^{n}\rangle$, are presented in the Supplement sup .
+Figure 3: Quench dynamics of nonintegrable Ising models on (a) one-dimensional, (b) square, and (c) cubic lattices.
 Shown is the expectation value of the total magnetization in the $z$-direction.
-The initial state is the pure product state (9) polarized along the same direction.
+The initial state is the pure product state (9) polarized along the same $z$ direction.
 Blue solid line is the result of the recursion method.
@@
 Points are the results of the sparse Pauli dynamics method Begušić and Chan (2025).
+Insets show explicitly computed Lanczos coefficients (points) and their extrapolation according to eqs.
+(4), (5) (black lines).
+The results of the recursion method are benchmarked by the results of exact diagonalization performed with QuSpin Weinberg and Bukov (2017) and time-dependent variational principle empowered by matrix product states performed with ITensor Fishman et al.
+(2022a, b) (see the Supplement sup for details).
+In one dimension these numerical methods are essentially exact up to sufficiently large times, therefore we regard their outcome as actual $\langle A_{t}\rangle$.
+One can see from Fig.
+1 that the recursion method excellently reproduces the actual $\langle A_{t}\rangle$ initially, but then abruptly breaks down.
+We observe that this breakdown is associated with a dynamical quantum phase transition (DQPT) Pollmann et al.
+(2010); Heyl et al.
+(2013); Karrasch and Schuricht (2013); Heyl (2018) – a phenomenon by which the Loschmidt rate function $\lambda(t)=N^{-1}\,\log\big{|}\langle\Psi_{0}|e^{-i\,t\,H}|\Psi_{0}\rangle\big{|}^{2}$ features singularities in the thermodynamic limit.111 There is an ambiguity in the meaning of the term “dynamical (quantum) phase transition.” Historically, it was introduced in a quite different context, referring to a transition between distinct phases of matter that unfolds over time due to variations in the Hamiltonian Schützhold et al.
+(2006); Eckstein et al.
+(2009); Uhlmann et al.
+(2010); Garrahan and Lesanovsky (2010); Diehl et al.
+(2010); Sciolla and Biroli (2010) – a definition more closely aligned with the conventional understanding of phase transitions.
+In contrast, the definition we adopt was introduced later Heyl et al.
+(2013) and is based on a formal analogy between the Loschmidt rate and the free energy.
+The breakdown time $t_{*}$ is very close to the time $t_{c}$ of one of the singularities (specifically, the second singularity), see Fig.
+1.
+More precisely, the breakdown time $t_{*}$ grows with the number $n_{\rm max}$ of available explicit Lanczos coefficient, but this growth saturates at $t_{c}$, as shown in Fig.
+2.
+This strengthens our belief that the observed association between the breakdown of the recursion method and the DQPT is not accidental.
+Remarkably, while the recursion method breaks down at a finite time $t_{*}$, at large times it seems to typically converge to the actual equilibrium value of the observable, as illustrated in Fig.
+3 (a).
+Note that the one-dimensional transverse-field Ising model is known for its unusually slow thermalization Bañuls et al.
+(2011); Lin and Motrunich (2017); Peng and Cui (2022); Pathak (2024) which is seen in Fig.
+3 (a).
 Quench in higher dimensions.
@@
 where ${\mathbf{i}}$ and ${\mathbf{j}}$ enumerate lattice sites and the first sum runs over pairs of neighbouring sites.
-We compute $22$ and $12$ nested commutators, respectively.
-Some results are presented in Fig.
+The observable is the same as in the previous case – the total magnetization in the $z$-direction, $A=\sum_{{\mathbf{j}}}\sigma^{z}_{{\mathbf{j}}}$.
+We compute $23$ and $12$ nested commutators, respectively.
+Time evolution of the magnetization is presented in Fig.
 3 (b),(c).
-They are compared to the outcomes of the sparse Pauli dynamics method Begušić and Chan (2024); Begušić .
+The results of the recursion method are compared to the outcomes of the sparse Pauli dynamics method Begušić and Chan (2025); Beg .
 One can see that the agreement is good initially but abruptly worsens at some threshold time.
@@
 However, we are not able to compute the critical time for this DQPT for two- and three-dimensional models.
-Conclusions.
+Conclusions and discussion.
 In this work, we have explored the applicability of the recursion method to computing the time evolution of observables relaxing from out-of-equilibrium initial states.
@@
 In one dimension, benchmarking is straightforward due to the availability of essentially exact numerical simulations.
-By comparing the recursion method to these simulations, we observed that it yields accurate results up to a certain threshold time, beyond which it abruptly breaks down.
-Notably, we found that this breakdown time closely coincides with the critical time of a dynamical quantum phase transition (DQPT), suggesting that DQPT fundamentally limits the applicability of the recursion method, at least in one dimension.
+By comparing the recursion method to these simulations, we have observed that it yields accurate results up to a certain threshold time, beyond which it abruptly breaks down.
+The breakdown indicates that the expansion (6) of a Heisenberg observable in the Lanczos basis has a finite convergence domain, at least for one-dimensional systems.
+This fact may limit the scope of concepts and techniques based on this expansion Parker et al.
+(2019); Chu et al.
+(2024); Loizeau et al.
+(2025c).
+Notably, the breakdown time turns out to closely coincide with the critical time of a dynamical quantum phase transition (DQPT), suggesting that DQPT fundamentally limits the applicability of the recursion method, at least in one dimension.
 The underlying mechanism linking DQPT to the breakdown of the recursion method remains entirely unknown and represents a pressing open question.
@@
 In contrast to the one-dimensional case, benchmark data in two and three dimensions is scarce.
-We have compared our results with those obtained from the sparse Pauli dynamics method Begušić and Chan (2024); Begušić and found that the two approaches diverge abruptly after a certain threshold time, similar to what is observed in one dimension.
+We have compared our results to those obtained from the sparse Pauli dynamics method Begušić and Chan (2025); Beg and found that the two approaches diverge abruptly after a certain threshold time, similar to what is observed in one dimension.
 However, due to the lack of available data on DQPTs in the higher-dimensional models we considered, we cannot unambiguously attribute this divergence to DQPTs.
-Determining whether DQPTs fundamentally constrain the recursion method’s convergence in higher dimensions remains an open question.
+Determining whether DQPTs fundamentally constrain the recursion method’s convergence in higher dimensions remains an important open question.
 Finally, in one dimension, we have observed that at long times, the recursion method’s outcome coincides with the time-averaged value of the observable, which, in thermalizing systems, is equal to the equilibrium value.
@@
 While this manuscript has been in preparation, a closely related study has appeared Loizeau et al.
-(2025b).
-The latter paper addresses quench dynamics in one-dimensional spin systems by means of the recursion method.
-It does not report neither the finite convergence domain of the method nor an association with the DQPT.
+(2025c).
+The latter paper also addresses quench dynamics in one-dimensional spin systems by means of the recursion method.
+It does not report the finite convergence domain of the method or any association with the DQPT.
 Acknowledgements.
@@
 We sincerely thank Alexey Rubtsov for pointing to the dynamical phase transition as a potential cause of the recursion method breakdown.
-We are grateful to Tomislav Begušić and Garnet Kin-Lic Chan for discussing their paper Begušić and Chan (2024), and particularly for providing unpublished data Begušić on the 2D Ising model.
-Numerous discussions with Alexander Teretenkov and Nikolay Il’in are acknowledged.
+We are grateful to Tomislav Begušić and Garnet Kin-Lic Chan for discussing their paper Begušić and Chan (2025) and our manuscript, and particularly for providing unpublished data Beg on the 2D Ising model.
+We thank Berislav Buča and Nicolas Loizeau for useful remarks and discussions.
+Finally, numerous discussions with Alexander Teretenkov and Nikolay Il’in are gratefully acknowledged.
 This work was supported by the Russian Science Foundation under grant № 24-22-00331, https://rscf.ru/en/project/24-22-00331/
+Appendix A Supplementary material
+Appendix B S1. Details on $\phi^{n}_{t}$ and $O^{n}_{t}$
+For a Hermitian operator $A$, the operator ${\mathcal{L}}^{n}A$ is Hermitian for even $n$ and anti-Hermitian for odd $n$.
+Therefore $O^{n}$ is Hermitian (anti-hermitian) for even (odd) $n$, and $\phi^{n}_{t}$ are real (imaginary) for even (odd) $n$.
+This $(-i)^{n}\phi^{n}_{t}$ and $i^{n}\langle O^{n}\rangle$ are always real.
+We plot these quantities in Fig.
+S1 and Fig.
+S2, respectively.
+Figure S1: Functions $(-i)^{n}\phi^{n}_{t}$ for $n=0,1,\dots,n_{\rm max}$.
+Shown are plots for (a) one-dimensional, (b) two-dimensional and (c) three-dimensional Ising models considered in the main text.
+As noted in the main text, $i^{n}\langle O^{n}\rangle$ do not exhibit any simple structure, as can be seen in Fig.
+S2.
+Importantly, available data show that convergence of $\big{|}\langle O^{n}\rangle\big{|}$ is either absent or, at least, rather slow.
+In fact, we have no reasons to expect fast convergence.
+Indeed, a straightforward sum rule for these coefficients,
+$\sum_{n=0}^{\infty}\big{|}\langle O^{n}\rangle\big{|}^{2}=\sum_{n=0}^{\infty}\left(N\,2^{N}\right)^{2}\,(\rho_{0}|O^{n})(O^{n}|\rho_{0})=\left(N\,2^{N}\right)^{2}\,(\rho_{0}|\rho_{0})=N\,2^{N}\operatorname{\mbox{tr}}\rho_{0}^{2}\geq N$ (S1)
+diverges in the thermodynamic limit of $N\rightarrow\infty$.
+This implies that $\big{|}\langle O^{n}\rangle\big{|}$ can not converge to zero faster than $1/\sqrt{n}$.
+It remains an interesting open question whether this slow (or absent) convergence is responsible for the breakdown of the recursion method.
+Figure S2: Values of $i^{n}\langle O^{n}\rangle$ for the one-dimensional (left column) and two-dimensional (right column) Ising models considered in the main text.
+Polarizations of the initial state are indicated in each plot.
+Appendix C S2. Numerics in the one-dimensional case
+For a one-dimensional model (7), we benchmark the time evolution of the observable by numerical methods applied to finite systems.
+Specifically, we perform exact diagonalization for up to $N=14$ spins with QuSpin Weinberg and Bukov (2017) and matrix product state-empowered time-dependent variational (TDVP) computation for up to $N=40$ spins.
+All numerical results coincide without visible discrepancies up to at least $t=3.5$, thus covering the time span of Fig.
+1.
+For larger times shown in Fig.
+3(a) we use the TDVP result for $N=40$.
+Appendix D S3. Moments and Lanczos coefficients
+Below we provide values of moments $\mu_{2n}=(A|{\mathcal{L}}^{2n}A)/(A|A)$ and Lanczos coefficients $b_{n}$ for models considered in the main text.
+An algorithm to obtain the latter from the former can be found in Dupuis (1967); Viswanath and Müller (2008).
+Note that, for technical reasons, in the 3D case we compute moments and Lanczos coefficients not for a translation-invariant observable (8), as for 1D and 2D cases, but for its local counterpart $\sigma^{z}_{{\mathbf{j}}}$.
+While moments and Lanczos coefficients of these two observables are different, their time dependent expectation values are equal up to an overall factor $N$ as long as the initial state is translation-invariant, thanks to the translation invariance of the Hamiltonian.
+$n$ $\mu_{2n}$ $b_{n}$
+1 12 3.4641
+2 480 5.2915
+3 25984 4.4934
+4 1694208 6.0281
+5 127258624 5.6190
+6 10783342592 6.9438
+7 1019673509888 6.6972
+8 107060323680256 8.2286
+9 12469034121428992 7.9953
+10 1613762515315982336 9.2837
+11 232736621970029805568 8.9148
+12 37468646645944075419648 10.2773
+13 6726394693098760675786752 9.8677
+14 1340433610285320688734568448 11.1094
+15 294391435294882867890188451840 10.7766
+16 70659698635880317667608199430144 12.1889
+17 18383424921170050832257258778263552 11.7912
+18 5147501922089241875609891362536685568 12.9532
+19 1542402400283276113669614948668265201664 12.7320
+20 492432525422001580271341582388177351475200 13.9288
+21 166985805198508106978965686562725967100379136 13.5778
+22 60011852488095956067395190665735924993040056320 14.6867
+23 22821321744324019152247154618657381361730948431872 14.4673
+24 9172207374041781749043802673339425314176975748726784 15.5798
+25 3892052853514868229060585022861205953766446860739805184 15.3892
+26 1741736488569588529172771998235278437950082051065922977792 16.4064
+27 821019133269536743838893539994338673397100678325485105577984 16.1096
+28 407086843763398331397385204208407255958886525287040747053252608 17.3132
+29 211990829889808789195437699387039831255135134538810758197371994112 17.0171
+30 115754651941408887232253140259122456149371760683740727687331373383680 18.0978
+31 66166596237703872256126915723522220984465097504279554678985792092635136 17.7211
+32 39530072936257449901374525090449349559603851447100658065034700516313006080 18.8846
+33 24646757783977889749525872721004379593240131044109640793744075941140506869760 18.6810
+34 16015858109930051337482877843244494350822380543617937802711923518036880757096448 19.6477
+35 10833753072406488462262074132859724923839003624262790509360495364480719760857759744 19.4833
+36 7620581541304124813656489403945273131158245029661872172588240591095613039962778238976 20.2878
+37 5568944585532111508410499104658395233823379770020424361966461875062472358794201533513728 20.3901
+38 4224439735905136953825989317559917466695928508761395026089245710815314351095999194227277824 21.0909
+39 3323843247775625289903213303600983380005695346567953993918161657219885198513483978643486015488 21.2234
+40 2710613205308213119249609362155540213052003237873428947613674131447949025132969096999139593420800 21.7286
+41 2289489892850970408834540810440144834655617422363836454740677715045134451758621541567258816170426368 21.9363
+42 2001453495607268189063620952290269505902583051218882018390254310106562921877530480766976590064753049600 22.6224
+43 1809588827470560013981821635403092794798688775869840545515811131487657718540062168703067251339768165826560 22.7147
+Table S1: Moments and Lanczos coefficients for the 1D Ising model (7) at $h_{x}=h_{z}=1$ and observable (8).
+Moments for $n=44,\dots,48$ are too large to fit the page.
+They can be obtained from the authors upon a request.
+$n$ $\mu_{2n}$ $b_{n}$
+1 16 4.0000
+2 896 6.3246
+3 71680 5.7966
+4 7307264 7.9714
+5 908099584 7.7771
+6 134882656256 9.8932
+7 23708042264576 9.8014
+8 4897950853496832 11.7753
+9 1181552437773729792 11.8055
+10 330379916698579894272 13.6724
+11 106222899499818563928064 13.7546
+12 38952574279318656784531456 15.6436
+13 16166117065306264371967557632 15.7188
+14 7539715274735761761366362292224 17.3406
+15 3926354218204565053083857654382592 17.3814
+16 2269213981250751881428716107876794368 19.0685
+17 1446793491966272434190827088127549505536 19.2419
+18 1011459407804179014361425714300323771711488 20.8675
+19 770720014993499216420856132351984701670424576 21.0498
+20 636516730341979806791563685949386538476858507264 22.6354
+21 566952387953410930810990384866974059612775307542528 22.8705
+22 542434116478048005253046423594975291775134675528318976 24.4221
+23 555699452970039450718350263160732485341995086705577689088 24.6588
+Table S2: Moments and Lanczos coefficients for the observable (8) and the Ising model on the square lattice, eq. (10), with $h_{z}=1$.
+$n$ $\mu_{2n}$ $b_{n}$
+1 24 4.8990
+2 1728 6.9282
+3 191232 7.6158
+4 28290048 9.3452
+5 5274918912 10.0996
+6 1206758473728 12.1273
+7 334929105125376 13.1690
+8 112136926246207488 14.8447
+9 44994911223284760576 15.9327
+10 21426816223192685740032 17.6780
+11 11966247190054308660903936 18.8192
+12 7744668425221733746879758336 20.5258
+Table S3: Moments and Lanczos coefficients for the observable $\sigma^{z}_{{\mathbf{j}}}$ and the Ising model on the cubic lattice, eq. (10), with $h_{z}=1$.

### v3

@@
-Quench dynamics via recursion method and dynamical quantum phase transitions
+Quench dynamics via recursion method
 Abstract
@@
 We apply the method to spin systems in one, two, and three spatial dimensions.
-In one dimension we find an evidence that the convergence of the method is limited in time by the dynamical quantum phase transition.
+In one dimension its usefulness is limited: although it remains accurate at moderate times, it eventually experiences an abrupt breakdown and subsequently yields results that deviate strongly from the true dynamics.
+In contrast, in two and three dimensions the method proves far more effective, providing a reasonably accurate description of the evolution across all time scales – from the initial transient regime to thermalization.
 Introduction.
@@
 Addressing the later in systems of dimension higher than one is known to be highly challenging, with limited body of results obtained to date Schmitt and Heyl (2020); Li et al.
-(2020); Dziarmaga (2022); Bernier and Agarwal (2025); Begušić and Chan (2025); Loizeau et al.
-(2025a, b).
+(2020); Dziarmaga (2022); Bernier and Agarwal (2025); Sinibaldi et al.
+(2024); Gravina et al.
+(2025); Pavešić et al.
+(2025); Chen et al.
+(2025); Begušić and Chan (2025); Loizeau et al.
+(2025a, b); Rudolph et al.
+(2025).
 Given the success of the recursion method in computing two-dimensional autocorrelation function, it seems tempting to apply it to quench dynamics.
@@
 The results are benchmarked by essentially exact numerical simulations.
-We demonstrate that the recursion method has a convergence domain limited in time by the dynamical quantum phase transition.
 Then results for the transverse-field Ising model on the square and cubic lattices are presented.
 We conclude by discussing implications of our results and open questions raised.
-[…]
-${\mathcal{L}}=\begin{pmatrix}0&b_{1}&0&0&\dots\\[3.00003pt] b_{1}&0&b_{2}&0&\dots\\[3.00003pt] 0&b_{2}&0&b_{3}&\dots\\[3.00003pt] 0&0&b_{3}&0&\dots\\[3.00003pt] \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}.$ (1)
+Recursion method for quench dynamics.
+Consider an observable given by a self-adjoint Shrödinger operator $A$.
+The same observable in the Heisenberg representation reads $A_{t}=e^{itH}A\,e^{-itH}$, where $H$ is the Hamiltonian of the system.
+The Heisenberg operator $A_{t}$ evolves within the Krylov space spanned by operators $A,{\mathcal{L}}A,{\mathcal{L}}^{2}A,\dots$, where ${\mathcal{L}}\equiv[H,\bullet]$ is the superoperator referred to as Liouvillian.
+$A_{t}$ satisfies the Heisenberg equation of motion $\partial_{t}A_{t}=i\,{\mathcal{L}}A_{t}$.
+Within the recursion method Viswanath and Müller (2008), this equation is addressed in the orthonormal Lanczos basis $\{O^{0},O^{1},O^{2},\dots\}$ constructed as follows: $O^{0}=A/\|A\|$, $A^{1}={\mathcal{L}}O^{0}$, $b_{1}=\|A^{1}\|$, $O^{1}=b_{1}^{-1}\,A^{1}$ and $A^{n}={\cal L}\,O^{n-1}-\,b_{n-1}O^{n-2}$, $b_{n}=\|A^{n}\|$, $O^{n}=b_{n}^{-1}A^{n}$ for $n\geq 2$.
+Here a scalar product in the space of operators is introduced according to $\big(A|B\big)\equiv{\cal N}\,\operatorname{\mbox{tr}}\big(A^{\dagger}B\big)$, where $\cal N$} is a normalization constant.
+For systems of $N$ spins $1/2$ and translation-invariant observables considered here it is convenient to choose ${\cal N}=N^{-1}\,2^{-N}$.
+The scalar product entails the norm $\|A\|=\sqrt{(A|A)}$.
+The coefficients $b_{n}$ are referred to as Lanczos coefficients.
+In the Lanczos basis the Liouvillian has a tridiagonal form,
+${\mathcal{L}}=\begin{pmatrix}0&b_{1}&0&0&\dots\\[3.00003pt] b_{1}&0&b_{2}&0&\dots\\[3.00003pt] 0&b_{2}&0&b_{3}&\dots\\[3.00003pt] 0&0&b_{3}&0&\dots\\[3.00003pt] \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}$.
+(1)
 The solution of the Heisenberg equation for the observable $A_{t}$ can be expressed as
@@
 (2025), we employ the following extrapolation formula:
-$b_{n}$ $\simeq\frac{\alpha\,n}{\log n}+\gamma+\frac{(-1)^{n}\alpha}{\log n+c}$ ${\rm for}~{}~{}1{\rm D},$ (4)
-$b_{n}$ $\simeq\alpha\,n+\gamma+\frac{(-1)^{n}\alpha}{\log n+c}$ ${\rm for}~{}~{}2,3{\rm D}.$ (5)
+$b_{n}$ $\simeq\frac{\alpha\,n}{\log n}+\gamma+\frac{(-1)^{n}\tilde{\alpha}}{\log n+c}$ ${\rm for}~~1{\rm D},$ (4)
+$b_{n}$ $\simeq\alpha\,n+\gamma+\frac{(-1)^{n}\tilde{\alpha}}{\log n+c}$ ${\rm for}~~2,3{\rm D}.$ (5)
 The last alternating term is required whenever $\varphi^{0}_{t}$ relaxes to a nonzero value at $t\rightarrow\infty$ Gamayun et al.
@@
 $\langle O^{n}\rangle$ is also computed symbolically.
+Final numerical values of $b_{n}$ are obtained with the use of rational arithmetic, which avoids the numerical instability of the Lanczos algorithm Eckseler et al.
+(2025).
 As for functions $\phi^{n}_{t}$, they are obtained by numerically solving coupled linear differential equations (3) with the tridiagonal matrix $\mathcal{L}$} truncated at some large $k$ (we use $k=1000$ in the examples below).
@@
 Below we explore it in specific models.
-Figure 1: Time evolution of the (a) Loschmidt rate and (b)–(d) total polarization along $z$ direction in the 1D Ising model (7) with $h_{x}=h_{z}=1$.
+Figure 1: Time evolution of the total polarization along $z$ direction in the 1D Ising model (7) with $h_{x}=h_{z}=1$.
 The initial state is a product translation-invariant state (9) with the polarization $(p_{x},p_{y},p_{z})$ indicated on each plot.
@@
 Blue line - results of the recursion method.
-The Loschmidt rate features multiple cusp singularities manifesting dynamical quantum phase transitions.
-The results of the recursion method depart from the exact numerical results at time $t_{*}$ close to the time $t_{c}$ of the second cusp (red dashed vertical line).
-Figure 2: Scaling of the estimated breakdown time $t_{*}$ with the number $n_{\rm max}$ of explicitly know Lanczos coefficients for various initial polarizations.
-$t_{*}$ is defined as the time when the recursion method result departs from the essentially exact numerical result by 0.005.
-The dashed horizontal line marks the time $t_{c}$ of the dynamical quantum phase transition, see Fig.
-1 (a).
-Clearly, $t_{*}$ exhibits a tendency to saturate around $t_{c}$.
+The recursion method reproduces initial evolution but abruptly departs from the exact numerical result at the breakdown time $t_{*}$ (vertical dashed line).
+This time appears to be independent of the initial state.
 Quench in one dimension.  We focus on a paradigmatic nonintegrable one-dimensional spin-$1/2$ chain – the mixed-field Ising model. Its Hamiltonian reads
@@
 One can see from Fig.
-1 that the recursion method excellently reproduces the actual $\langle A_{t}\rangle$ initially, but then abruptly breaks down.
-We observe that this breakdown is associated with a dynamical quantum phase transition (DQPT) Pollmann et al.
-(2010); Heyl et al.
-(2013); Karrasch and Schuricht (2013); Heyl (2018) – a phenomenon by which the Loschmidt rate function $\lambda(t)=N^{-1}\,\log\big{|}\langle\Psi_{0}|e^{-i\,t\,H}|\Psi_{0}\rangle\big{|}^{2}$ features singularities in the thermodynamic limit.111 There is an ambiguity in the meaning of the term “dynamical (quantum) phase transition.” Historically, it was introduced in a quite different context, referring to a transition between distinct phases of matter that unfolds over time due to variations in the Hamiltonian Schützhold et al.
-(2006); Eckstein et al.
-(2009); Uhlmann et al.
-(2010); Garrahan and Lesanovsky (2010); Diehl et al.
-(2010); Sciolla and Biroli (2010) – a definition more closely aligned with the conventional understanding of phase transitions.
-In contrast, the definition we adopt was introduced later Heyl et al.
-(2013) and is based on a formal analogy between the Loschmidt rate and the free energy.
-The breakdown time $t_{*}$ is very close to the time $t_{c}$ of one of the singularities (specifically, the second singularity), see Fig.
-1.
-More precisely, the breakdown time $t_{*}$ grows with the number $n_{\rm max}$ of available explicit Lanczos coefficient, but this growth saturates at $t_{c}$, as shown in Fig.
-2.
-This strengthens our belief that the observed association between the breakdown of the recursion method and the DQPT is not accidental.
-Remarkably, while the recursion method breaks down at a finite time $t_{*}$, at large times it seems to typically converge to the actual equilibrium value of the observable, as illustrated in Fig.
+1 that the recursion method excellently reproduces the actual $\langle A_{t}\rangle$ initially, but then abruptly breaks down at the breakdown time $t=t_{*}$.
+After this breakdown, the recursion method result strongly deviates from the actual value of the observable.
+We note that the breakdown time appears to be independent of the initial state.
+We attribute this breakdown of the recursion method to the slow thermalization typical of one-dimensional systems.
+In particular, the mixed-field Ising model is well known for its anomalous thermalization behavior Bañuls et al.
+(2011); Lin and Motrunich (2017); Kormos et al.
+(2017); Peng and Cui (2022); Pathak (2024).
+In the previous version of this manuscript Shirokov et al.
+(2025), we linked the breakdown of the recursion method to a dynamical quantum phase transition (DQPT), motivated by an apparent coincidence between the breakdown time and the DQPT time.
+A more careful analysis shows that this coincidence is not universal and does not hold across most of the parameter space.
+We therefore retract our earlier claim that a DQPT is the cause of, or a universal precursor to, the breakdown of the recursion method in one dimension.
+Additional data and discussion on this point are provided in the Supplement.
+Remarkably, at large times the recursion method result typically converges to the actual equilibrium value of the observable, as illustrated in Fig.
 2 (a).
-Note that the one-dimensional transverse-field Ising model is known for its unusually slow thermalization Bañuls et al.
-(2011); Lin and Motrunich (2017); Peng and Cui (2022); Pathak (2024) which is seen in Fig.
-3 (a).
+This convergence may indicate that the terms discarded by the truncation in Eq.
+(2), while significantly affecting the transient dynamics, do not contribute much in the limit of $t\rightarrow\infty$.
 Quench in higher dimensions.
@@
 The observable is the same as in the previous case – the total magnetization in the $z$-direction, $A=\sum_{{\mathbf{j}}}\sigma^{z}_{{\mathbf{j}}}$.
-We compute $23$ and $12$ nested commutators, respectively.
+We compute $23$ and $17$ nested commutators, respectively.
 Time evolution of the magnetization is presented in Fig.
 2 (b),(c).
-The results of the recursion method are compared to the outcomes of the sparse Pauli dynamics method Begušić and Chan (2025); Beg .
-One can see that the agreement is good initially but abruptly worsens at some threshold time.
-We are not able to decisively resolve the origin of this discrepancy at the moment.
-Of course, the DQPT is again the main suspect.
-DQPT is known to occur in the 2D Ising model, although with a softer logarithmic singularity Heyl (2018).
-However, we are not able to compute the critical time for this DQPT for two- and three-dimensional models.
+The results of the recursion method are compared to the outcomes of the sparse Pauli dynamics (SPD) method Begušić and Chan (2025); Beg .
+Up to the maximal time accessible to the SPD, the agreement with our results is excellent at early times and remains reasonably good thereafter.
+In contrast to the SPD, however, the recursion method yields predictions all the way to equilibration.
 Conclusions and discussion.
@@
 By comparing the recursion method to these simulations, we have observed that it yields accurate results up to a certain threshold time, beyond which it abruptly breaks down.
-The breakdown indicates that the expansion (6) of a Heisenberg observable in the Lanczos basis has a finite convergence domain, at least for one-dimensional systems.
-This fact may limit the scope of concepts and techniques based on this expansion Parker et al.
+Despite this breakdown, the recursion method result eventually equilibrates close to the actual equilibrium value.
+The breakdown indicates the slow convergence of the expansion (6) of a Heisenberg observable in the Lanczos basis.
+This slow convergence may limit the scope of concepts and techniques Parker et al.
 (2019); Chu et al.
 (2024); Loizeau et al.
-(2025c).
-Notably, the breakdown time turns out to closely coincide with the critical time of a dynamical quantum phase transition (DQPT), suggesting that DQPT fundamentally limits the applicability of the recursion method, at least in one dimension.
-The underlying mechanism linking DQPT to the breakdown of the recursion method remains entirely unknown and represents a pressing open question.
-The potential subtlety of this mechanism is highlighted by our observation that the breakdown time coincides not with the first, but with the second (in time) dynamical critical point, implying that not every DQPT necessarily disrupts the recursion method.
-In contrast to the one-dimensional case, benchmark data in two and three dimensions is scarce.
-We have compared our results to those obtained from the sparse Pauli dynamics method Begušić and Chan (2025); Beg and found that the two approaches diverge abruptly after a certain threshold time, similar to what is observed in one dimension.
-However, due to the lack of available data on DQPTs in the higher-dimensional models we considered, we cannot unambiguously attribute this divergence to DQPTs.
-Determining whether DQPTs fundamentally constrain the recursion method’s convergence in higher dimensions remains an important open question.
-Finally, in one dimension, we have observed that at long times, the recursion method’s outcome coincides with the time-averaged value of the observable, which, in thermalizing systems, is equal to the equilibrium value.
-If this observation proves to be correct and generic, it would imply an unusual scenario: the recursion method remains valid at both short and long times but fails at intermediate times.
-Investigating this phenomenon further presents another promising direction for future research.
+(2025c); Alishahiha and Vasli (2025); Rabinovici et al.
+(2025) built on this expansion in one dimension.
+In contrast to the one-dimensional case, benchmark data in two – and especially three – dimensions remain scarce.
+We have compared our results with those obtained using the SPD method Begušić and Chan (2025); Beg and find agreement ranging from excellent to reasonable up to the maximal times accessible to SPD.
+Importantly, the recursion method is essentially unrestricted in time and continues to yield reasonable results all the way to equilibration.
 Note added.
@@
 (2025c).
-The latter paper also addresses quench dynamics in one-dimensional spin systems by means of the recursion method.
-It does not report the finite convergence domain of the method or any association with the DQPT.
+The latter paper addresses quench dynamics in one-dimensional spin systems by means of the recursion method.
 Acknowledgements.
 Acknowledgments.
-We sincerely thank Alexey Rubtsov for pointing to the dynamical phase transition as a potential cause of the recursion method breakdown.
 We are grateful to Tomislav Begušić and Garnet Kin-Lic Chan for discussing their paper Begušić and Chan (2025) and our manuscript, and particularly for providing unpublished data Beg on the 2D Ising model.
@@
 For larger times shown in Fig.
-3(a) we use the TDVP result for $N=40$.
+2(a) we use the TDVP result for $N=40$.
 Appendix D S3. Moments and Lanczos coefficients
@@
 While moments and Lanczos coefficients of these two observables are different, their time dependent expectation values are equal up to an overall factor $N$ as long as the initial state is translation-invariant, thanks to the translation invariance of the Hamiltonian.
+Note that in the previous version of the manuscript Shirokov et al.
+(2025), for the 3D Ising model, we provided 12 moments and Lanczos coefficients for the non-translation-invariant observable $\sigma^{z}_{j}$.
+Here we provide 17 moments and Lanczos coefficients for the respective translation-invariant observable.
+While the values are different in the two cases, the output of the recursion method is consistent whenever the initial state is translation-invariant.
 $n$ $\mu_{2n}$ $b_{n}$
@@
 $n$ $\mu_{2n}$ $b_{n}$
-1 24 4.8990
-2 1728 6.9282
-3 191232 7.6158
-4 28290048 9.3452
-5 5274918912 10.0996
-6 1206758473728 12.1273
-7 334929105125376 13.1690
-8 112136926246207488 14.8447
-9 44994911223284760576 15.9327
-10 21426816223192685740032 17.6780
-11 11966247190054308660903936 18.8192
-12 7744668425221733746879758336 20.5258
+1 24 4.89898
+2 1920 7.48331
+3 230400 7.55929
+4 36618240 9.98284
+5 7313326080 10.1916
+6 1790621122560 12.7955
+7 530848042450944 13.1431
+8 188884569821282304 15.4032
+9 79892590021627084800 15.8832
+10 39721461646806598287360 18.2052
+11 22947716425271653039079424 18.7155
+12 15244534918363394241158184960 20.9982
+13 11543604835131049187075722051584 21.5619
+14 9891579029630566720438258058133504 23.8488
+15 9532138756054379075322538513856987136 24.4597
+16 10273436716844388765436471742460667625472 26.751
+17 12321137907981191488465349016456502941057024 27.4636
 Table S3: Moments and Lanczos coefficients for the observable $\sigma^{z}_{{\mathbf{j}}}$ and the Ising model on the cubic lattice, eq. (10), with $h_{z}=1$.
+Appendix E S4. Recursion method breakdown and dynamical quantum phase transition in one dimension
+Figure S3: Time evolution of the total polarization along $z$ direction in the 1D Ising model (7) for various values of $h_{z}$ and $h_{x}$.
+The initial state is a product translation-invariant state (9) polarized in the $z$-direction.
+Orange line – essentially exact results of the numerical time evolution (see text for details).
+Blue line – results of the recursion method.
+Pink line – the rate function $\lambda(t)$ whose cusps mark the dynamical quantum phase transitions.
+In the previous version of this manuscript Shirokov et al.
+(2025), we have attributed the breakdown of the recursion method to a dynamical quantum phase transition (DQPT) Pollmann et al.
+(2010); Heyl et al.
+(2013); Karrasch and Schuricht (2013); Heyl (2018) – a phenomenon by which the Loschmidt rate function $\lambda(t)=-N^{-1}\,\log\big|\langle\Psi_{0}|e^{-i\,t\,H}|\Psi_{0}\rangle\big|^{2}$ features singularities in the thermodynamic limit.
+[^1] There is an ambiguity in the meaning of the term “dynamical (quantum) phase transition.” Historically, it was introduced in a quite different context, referring to a transition between distinct phases of matter that unfolds over time due to variations in the Hamiltonian Schützhold et al.
+(2006); Eckstein et al.
+(2009); Uhlmann et al.
+(2010); Garrahan and Lesanovsky (2010); Diehl et al.
+(2010); Sciolla and Biroli (2010) – a definition more closely aligned with the conventional understanding of phase transitions.
+In contrast, the definition we adopt was introduced later Heyl et al.
+(2013) and is based on a formal analogy between the Loschmidt rate and the free energy.
+Here, we have examined this conjecture more thoroughly across a range of model parameters.
+The results, shown in Fig.
+S3, indicate that, in general, $t_{*}$ and $t_{c}$ do not coincide.
+The apparent agreement occurs, within uncertainties, only for $h_{x}\simeq h_{z}$.
+We therefore retract our earlier strong claim regarding a relation between the recursion method breakdown and the DQPT, leaving further investigation of this connection to future work.
+For the sake of this future work, we point to a paper Takahashi (2025) on DQPT and Lanczos method in the Schrödinger representation.